OpenStudy (anonymous):

A laptop computer was originally priced at $695.00. The price was reduced by 9%. What is the new price of the laptop?

OpenStudy (sleepyjess):

695*.09

OpenStudy (sleepyjess):

9% in decimal form is just .09, so all you have to do is multiply.

OpenStudy (anonymous):

MAY I ASK HOW?

OpenStudy (sleepyjess):

???

OpenStudy (anonymous):

62.55

OpenStudy (sleepyjess):

Ok, now subtract that from the original price

OpenStudy (anonymous):

632.45

OpenStudy (doc.brown):

Per-cent means "out of one hundred". So if you chop $695.00 into 100 pieces, you only want 9 of them.\[\frac{$695.00}{100}\times 9\]Sleepyjess did some of the work for you.\[695\times\frac{9}{100}\]\[695\times 0.09\]\[=62.55\]Remember, you want to reduce the price of the computer by 9%. \[$695.00-\left($695.00\times\frac{9}{100}\right)\]

OpenStudy (sleepyjess):

Awesome explanation @doc.brown